# Discrete Diffusion in Large Language and Multimodal Models: A Survey

### v3

@@ II.A Discrete Diffusion Model and Transition Matrix
    $[Q_{t}]_{ij}=\begin{cases}\frac{\exp\left(-\frac{4|i-j|^{2}}{(K-1)^{2}\beta_{t}}\right)}{\sum\limits_{n=-(K-1)}^{K-1}\exp\left(-\frac{4n^{2}}{(K-1)^{2}\beta_{t}}\right)}&\text{if }i\neq j\\ 1-\sum\limits_{l\neq i}[Q_{t}]_{il}&\text{if }i=j\end{cases}$
@@
    $[Q_{t}]_{ij}=\begin{cases}\frac{1}{K\beta_{t}}&0<|i-j|\leq v\\ 0&|i-j|>v\\ 1-\sum\limits_{l\neq i}[Q_{t}]_{il}&i=j\end{cases}$
@@
    $R_{ij}$ =$=\begin{cases}-\sum_{l\neq i}A_{il}&i=j\\ A_{ij}&i\neq j\end{cases}.$ (10)
@@ II.C Simplified Masked Diffusion Model
By defining the discrete time series as $10,\frac{1}{T},\ldots,1-\frac{1}{T},$ 1 and letting $T\to\infty$, both works [16, 17] extend the above loss Eq. 29 to the continuous-time setting:
@@ II.D Continuous Time Discrete Denoising Models
During inference, however, the exact simulation of the reverse CTMC can be computationally prohibitive. Instead, the tau-leaping algorithm [reference] approximates the reverse process by applying multiple transitions within a time interval $\tau$ simultaneously. For a current state $x_{t}$, the number of transitions to each $x^{\prime}$ during $[t-\tau,t]$ is modeled as:
@@ II.E.1 Training Loss
$\ell^{i}_{\text{score}}=\mathcal{D}\left(\left[\frac{p_{1|t}([y^{i},x_{1}^{\backslash i}]|x_{t})}{p_{1|t}(x_{1}|x_{t})}\right]_{y^{i}=1}^{V},\left[\frac{p^{\theta}_{1|t}([y^{i},x_{1}^{\backslash i}]|x_{t})}{p_{\theta}^{1|t}(x_{1}|x_{t})}\right]_{y^{i}=1}^{V}\right),$ (40)
where $V$ is the vocabulary size, $\mathcal{D}(\cdot,\cdot)$ denotes a divergence measure, $\omega(t)$ is the distribution used to sample time index $t$, and $h(x_{1}|x_{t})$ is a proposal distribution such as the ground-truth conditional distribution $p_{1|t}(x_{1}|x_{t})$. In the equation, $[y^{i},x^{\backslash i}]:=[x^{1},\ldots,x^{i-1},y^{i},x^{i+1},\ldots,x^{L}]$ is used to define a new sequence of the same length as $x_{1}$, where the token at position $i$ is replaced by $y^{i}$ and all other tokens remain identical to those in $x_{1}$. $x^{\backslash i}_{1}$ is used to indicate all tokens in $x_{1}$ except for the token at position $i$.
@@
$\mathcal{L}_{\mathrm{distrib}}(\theta)=\mathbb{E}_{\omega(t)p(x_{t})}\sum_{i=1}^{L}\mathbb{E}_{h(x_{1}^{\backslash i}|x_{t})}\ell^{i}_{\text{distrib}},$ (41)
$\ell^{i}_{\text{distrib}}=\mathbb{D}\left(p_{1|t}(x_{1}^{i}|x_{1}^{\backslash i},x_{t})\,\|\,p^{\theta}_{1|t}(x_{1}^{i}|x_{1}^{\backslash i},x_{t})\right),$ (42)
@@ II.F Discrete Flow Matching (DFM)
II.G Block Diffusion Models
@@
$-\log p_{\theta}(x)\leq\mathcal{L}_{\mathrm{BD}}(x;\theta):=\sum_{b=1}^{B}\mathcal{L}(x^{(b)},x^{(<b)};\theta),$ (64)
@@
III Large Discrete Diffusion Language and Multimodal Models
@@
III.A Discrete Diffusion Models around 1B
@@
III.B Large Diffusion Language Models
III.B.1 LLaDA
[6] introduces LLaDA, the first discrete diffusion-based alternative to autoregressive LLMs. Instead of predicting one token at a time in a left-to-right manner, LLaDA generates text by gradually denoising masked text: a forward process randomly masks tokens in the input, and a reverse process uses a Transformer to predict the masked tokens. This Transformer mask predictor has an architecture similar to standard LLMs, but notably omits the causal masking, allowing bidirectional context since the model sees the entire sequence during prediction. LLaDA follows a principled generative modeling approach by optimizing a variational likelihood bound (ELBO) [83] rather than the exact log-likelihood. This design leverages the full context of text and the theoretical grounding of diffusion models to potentially achieve LLM capabilities beyond the autoregressive paradigm.
LLaDA demonstrates that dLLMs can attain competitive performance with same-size LLMs. Specifically, LLaDA 8B achieved comparable zero-/few-shot accuracy to LLaMA-3 8B [84] on a suite of 15 standard tasks [6], despite not using an autoregressive decoder. After instruction-tuning, LLaDA shows strong ability to follow complex instructions in dialogues, comparable to other fine-tuned LLMs. Notably, due to its bidirectional generation, LLaDA overcomes certain challenges that stump AR models: for example, it experimentally breaks the ‘reversal curse’ [6] and can correctly complete a reversed-order poem task, surpassing even GPT-4 in such evaluation [83]. These results establish that dLLMs are a viable and promising alternative to LLMs.
III.B.2 DIFFUSION-LLMs
@@ III.B.3 DiffuGPT & DiffuLLaMA
III.B.4 DREAM
DREAM 7B [7] is one of the most powerful open-source dLLMs to date. The name ‘DREAM’ underscores its focus on diffusion for reasoning and modeling—it is specifically designed to excel at complex reasoning tasks while remaining efficient and scalable. DREAM 7B achieves performance on par with, or exceeding, autoregressive models of similar size (e.g. it matches LLaMA3 8B [84] and Qwen 2.5-7B [90] on many benchmarks). A key to DREAM’s success is an optimized training recipe distilled from extensive experiments at smaller scales. [7] carefully explores the design choices on a 1B model and identify two especially valuable components: (1) AR weight initialization and (2) context adaptive noise scheduling.
@@ III.B.4 DREAM
III.B.5 LLaDA 1.5
Another technique in VRPO is antithetic sampling for preference comparisons. In preference optimization (e.g. Direct Preference Optimization or DPO [92]), the model compares the log-likelihoods of a better output vs a worse output. VRPO uses the same random noise for both outputs when estimating their ELBO scores. By sharing diffusion timesteps and mask patterns between the winning and losing outputs, the random errors in their log-likelihood estimates tend to cancel out, greatly reducing variance in the difference. This is analogous to using the same random seed for paired evaluations to get a more reliable comparison. With these methods—increased samples, optimal allocation, and antithetic sampling—VRPO achieves a significant stability improvement in aligning the diffusion model. [43] applys VRPO to LLaDA and fine-tune it with a preference model (reward model) on reasoning, math, code, and some other general alignment tasks. The resulting LLaDA 1.5 shows a marked performance jump over the SFT-only LLaDA on all evaluated fronts. In summary, LLaDA 1.5 demonstrates that RLHF-style alignment is feasible for dLLMs and that addressing the ELBO variance issue is the key perspective.
III.B.6 TESS 2
@@
III.C Large Diffusion Multimodal Models
III.C.1 Dimple
@@
During inference, Dimple introduces a confident decoding strategy for efficiency: the model dynamically chooses how many tokens to fill in at each step based on model confidence (see Sec. V.A). Empirically, this reduces the number of iterations to about response $\frac{\text{response length}}{3}$. Dimple also re-implements an autoregressive prefilling trick: by filling in some tokens from the existing context, it speeds up inference by about 1.5$\times$–7$\times$ with minimal impact on quality. Under the same training budget and dataset as LLaVA-NEXT [97], Dimple-7B achieves higher aggregate scores on multimodal benchmarks than LLaVA-NEXT-7B. This result shows that with a proper hybrid training recipe, a discrete dMLLM can match or exceed strong autoregressive baselines on vision-language tasks.
III.C.2 LaViDa
@@ III.C.3 LLaDA-V
III.D Large Unified Model
III.D.1 MMaDA
@@ IV.A.1 Low Corpus Utilization
IV.A.2 Random Sampling of Time Index
@@ IV.B.1 BERT Initialization
IV.B.2 Autoregressive Model Initialization
@@
IV.B.3 Autoregressive-then-Diffusion Training
@@
This hybrid approach addresses several known limitations of pure diffusion training—such as low corpus utilization and length bias—by initializing the model with strong AR priors. With this hybrid training approach, Dimple achieves better alignment, improved instruction tuning, and enhanced robustness to response length variations.
IV.C Complementary Masking Technique
In [10], to ensure that all tokens participate in training, complementary masking constructs two complementary masked versions of each input sequence: $X_{t}$ and $X_{t}^{C}$. These versions have non-overlapping masked spans. For example, consider the sentence:
@@ IV.C Complementary Masking Technique
IV.D Masking Scheduling Technique
@@
IV.D.1 Uniform Masking Scheduling
@@
IV.D.2 Token-wise Masking Scheduling
Uniform masking scheduling uses the same scheduling for all tokens, which ignores the inherent variation in informativeness among different tokens. For example, different tokens carry different amounts of information, typically measured by entropy and dLLMs exhibit an easy-first decoding behavior—i.e., common, low-entropy tokens are easier to predict earlier. [35] introduces a token-wise masking schedule.
The schedule defines a custom corruption probability $\alpha_{t}^{i}$ for each token position $i$ at timestep $t$, determined by:
@@ IV.D.2 Token-wise Masking Scheduling
where
  • $H(x_{0}^{i})$ denotes the entropy of the $i$-th token, measuring its information content.
  • $S(t)$ is a sinusoidal scaling function that ensures zero informativeness contribution at $t=0$ and $t=T$.
  • $\lambda$ is a hyperparameter controlling the strength of entropy influence.
@@
IV.E Addressing Training-Testing Discrepancy

### v4

@@ II Mathematical Formulations
In this section, we discuss the mathematical formulation of the discrete diffusion.
@@ II.A Discrete Diffusion Model and Transition Matrix
    $Q_{t}]_{ij}=\begin{cases}\frac{\exp\left(-\frac{4|i-j|^{2}}{(K-1)^{2}\beta_{t}}\right)}{\sum\limits_{n=-(K-1)}^{K-1}\exp\left(-\frac{4n^{2}}{(K-1)^{2}\beta_{t}}\right)}&\text{if }i\neq j\\ 1-\sum\limits_{l\neq i}[Q_{t}]_{il}&\text{if }i=j\end{cases}$
@@
    $Q_{t}]_{ij}=\begin{cases}\frac{1}{K\beta_{t}}&0<|i-j|\leq v\\ 0&|i-j|>v\\ 1-\sum\limits_{l\neq i}[Q_{t}]_{il}&i=j\end{cases}$
@@
    $R_{ij}$ $=\begin{cases}-\sum_{l\neq i}A_{il}&i=j\\ A_{ij}&i\neq j\end{cases}.$ (10)
@@ II.C Simplified Masked Diffusion Model
By defining the discrete time series as $0,\frac{1}{T},\ldots,1-\frac{1}{T},1$ and letting $T\to\infty$, both works [19, 20] extend the above loss Eq. 29 to the continuous-time setting:
@@ II.D Continuous Time Discrete Denoising Models
During inference, however, the exact simulation of the reverse CTMC can be computationally prohibitive. Instead, the tau-leaping algorithm [reference] approximates the reverse process by applying multiple transitions within a time interval $\tau$ simultaneously. For a current state $x_{t}$, the number of transitions to each $x^{\prime}$ during $t-\tau,t]$ is modeled as:
@@ II.E.1 Training Loss
[…]
where $V$ is the vocabulary size, $\mathcal{D}(\cdot,\cdot)$ denotes a divergence measure, $\omega(t)$ is the distribution used to sample time index $t$, and $h(x_{1}|x_{t})$ is a proposal distribution such as the ground-truth conditional distribution $p_{1|t}(x_{1}|x_{t})$. In the equation, $y^{i},x^{\backslash i}]:=[x^{1},\ldots,x^{i-1},y^{i},x^{i+1},\ldots,x^{L}]$ is used to define a new sequence of the same length as $x_{1}$, where the token at position $i$ is replaced by $y^{i}$ and all other tokens remain identical to those in $x_{1}$. $x^{\backslash i}_{1}$ is used to indicate all tokens in $x_{1}$ except for the token at position $i$.
@@
[…]
@@ II.F Discrete Flow Matching (DFM)
III Modeling Language Diffusion
Beyond the mathematical formulation introduced in the previous section, in this section, we present several techniques used when modeling the language diffusion task, such as specialized neural network designs, and extra sub-tasks. These techniques are introduced to enhance the controllability and flexibility.
III.A Block Diffusion Models
@@
$\log p_{\theta}(x)\leq\mathcal{L}_{\mathrm{BD}}(x;\theta):=\sum_{b=1}^{B}\mathcal{L}(x^{(b)},x^{(<b)};\theta),$ (64)
@@
III.B Flexible-Length Masked Diffusion
Flexible-Length Masked Diffusion (FlexMDM) [39] extends masked diffusion models to handle variable-length sequences.At each step of the inference process, FlexMDM inserts additional mask tokens into the current response token sequence, thereby enabling variations in the response length.
In FlexMDM, each token may reside in one of three states: empty, masked, or ground truth (gt). In the forward process, a token first transitions from the ground truth state to the masked state, and finally to the empty state. Let $x_{1}=(x^{1}_{1},\ldots,x^{n}_{1})\sim p_{1}$ be a target sequence of length $n$. FlexMDM defines two smooth monotone schedules $\alpha,\beta:[0,1]\to[0,1]$ with boundary conditions $(\alpha_{0},\beta_{0})=(0,0)$ and $(\alpha_{1},\beta_{1})=(1,1)$. For each token $i$, FlexMDM independently sample an insertion time $T^{i}_{1}$ and an unmasking time $T^{i}_{2}$:
$T^{i}_{1}\sim\dot{\alpha}_{t}\,dt,\quad T^{i}_{2}\sim\mathbf{1}\{t\geq T^{i}_{1}\}\,\frac{\dot{\beta}_{t}}{1-\beta_{T^{i}_{1}}}\,dt,$
where $\dot{\alpha}_{t}$ and $\dot{\beta}_{t}$ are the derivatives of $\alpha_{t}$ and $\beta_{t}$. The token state $x^{i}_{t}$ at time $t\in[0,1]$ evolves as:
$x^{i}_{t}=\begin{cases}\emptyset,&0<t<T^{i}_{1},\\ m,&T^{i}_{1}\leq t<T^{i}_{2},\\ x^{i}_{1},&T^{i}_{2}\leq t\leq 1,\end{cases}$
where $m$ denotes the special mask symbol and $\emptyset$ represents token removal. The sequence $x_{t}$ is obtained by concatenating all non-empty $x^{i}_{t}$.
Thus, in FlexMDM, there are two complementary tasks:
  1. 1.
    Unmasking Task. Given a partially masked sequence $x_{t}$, the model predicts the ground-truth token for each masked position. This is identical to the original MDM formulation. Specifically, a diffusion language model $f_{\theta}$ is trained to approximate the posterior distribution over clean tokens.
  2. 2.
    Insertion Task. Beyond unmasking, FlexMDM introduces the additional task of inserting tokens into the sequence. For this purpose, an auxiliary network $g_{\theta}$ is trained to predict the expected number of mask tokens that should be inserted before each existing token. During inference, this prediction is combined with a Poisson distribution to determine the actual number of insertions:
    $k_{i}\sim\text{Poisson}\!\left(r\cdot g_{\theta}(x_{t},t)[i]\right),$
    where $r$ is a scaling factor. The sampled $k$ tokens are inserted as mask symbols $m$, thereby extending the sequence length.
At inference time, sequence generation alternates between the two tasks:
  1. 1.
    Insertion: For each position $i$, sample $k_{i}\sim\text{Poisson}(r\cdot g_{\theta}(x_{t},t)[i])$ and insert $k_{i}$ mask tokens into the sequence.
  2. 2.
    Unmasking: Apply $f_{\theta}$ to predict clean tokens at masked positions and replace them accordingly.
This insertion–unmasking cycle is repeated iteratively until the sequence converges to a fully unmasked form, thereby producing variable-length outputs while preserving the any-order property of diffusion-based inference.
III.C Partial Masking
In a standard Masked Diffusion Models, each token can only exist in one of two states: masked or unmasked. Partial Masking [40] scheme extends MDM by introducing intermediate states, achieved by decomposing each token into a sequence of sub-tokens.
Each original token $x^{i}_{0}\in\mathcal{X}$, where $\mathcal{X}=\{0,\ldots,C-1\}$, is mapped into a sub-token sequence $y^{i}_{0}=[y^{i,1}_{0},\ldots,y^{i,\ell}_{0}]\in\mathcal{Y}^{\ell}$ through an invertible base-$b$ encoding function $f$:
$f:\mathcal{X}\to\mathcal{Y}^{\ell},\quad y^{i}_{0}=f(x^{i}_{0}),$
where $\mathcal{Y}=\{0,\ldots,b-1\}$ and $b=\lceil\sqrt[\ell]{C}\rceil$. The inverse mapping $f^{-1}$ reconstructs tokens from sub-tokens, ensuring lossless transformation:
$x^{i}_{0}=f^{-1}(y^{i}_{0}).$
Consequently, the forward pass, backward pass, training, and inference of diffusion with partial masking all occur at the sub-token level.
III.D Discrete Diffusion with Optimal Transport Position Coupling
Standard discrete diffusion models relies on fixed token positions during generation, which prevents them from handling flexible-length or flexible-position text infilling. Discrete Diffusion with Optimal Transport Position Coupling (DDOT) [41] jointly denoises both token values and token positions, enabling dynamic sequence restructuring while preserving relative ordering.
At each timestep $t$, DDOT outputs both (i) the predicted token value distribution and (ii) the velocity of token positions:
$s_{\theta}(x_{t},t)\quad\text{and}\quad v_{\theta}(z_{t},t),$
where $x_{t}$ are token values and $z_{t}\in[-1,1]^{L}$ are continuous position variables. The extra velocity of token position is predicted by and additional linear head. The token positions are diffused in continuous space from an initial distribution $z_{T}\sim U(-1,1)^{L}$ to the ground-truth aligned positions $z_{0}$. The token denoising objective follows the score entropy loss from SEDD [47], while the position denoising is learned via a weighted mean-squared error loss:
$\mathcal{L}_{\text{pos}}(\theta)=\mathbb{E}_{(z,t)}\Big{[}Q_{t}(x_{t},y)\,\|v_{\theta}(z_{t},t)-(z_{0}-z_{T})\|^{2}\Big{]}.$
During sampling, DDOT alternates between token denoising and position refinement:
  1. 1.
    Predict token distributions $s_{\theta}(x_{t},t)$ and replace masks accordingly.
  2. 2.
    Predict position velocities $v_{\theta}(z_{t},t)$ and update $z_{t}$ with Euler steps.
@@ III.D Discrete Diffusion with Optimal Transport Position Coupling
IV Large Discrete Diffusion Language and Multimodal Models
@@
IV.A Discrete Diffusion Models around 1B
@@
IV.B Large Diffusion Language Models
IV.B.1 LLaDA Series
The LLaDA series represents the pioneering line of discrete diffusion-based alternatives to autoregressive LLMs. LLaDA [6] is the first diffusion language model designed to replace autoregressive decoding with a mask-predict-denoise paradigm. Instead of generating text token by token in a left-to-right manner, LLaDA gradually reconstructs masked text sequences through a forward masking and reverse denoising process. The mask predictor is implemented with a Transformer architecture without causal masking, enabling bidirectional context utilization. LLaDA follows a principled generative modeling approach by optimizing a variational likelihood bound (ELBO) [124] rather than the exact log-likelihood. This design leverages the full context of text and the theoretical grounding of diffusion models to potentially achieve LLM capabilities beyond the autoregressive paradigm.
Another technique in VRPO is antithetic sampling for preference comparisons. In preference optimization (e.g. Direct Preference Optimization or DPO [126]), the model compares the log-likelihoods of a better output vs a worse output. VRPO uses the same random noise for both outputs when estimating their ELBO scores. By sharing diffusion timesteps and mask patterns between the winning and losing outputs, the random errors in their log-likelihood estimates tend to cancel out, greatly reducing variance in the difference. This is analogous to using the same random seed for paired evaluations to get a more reliable comparison. With these methods—increased samples, optimal allocation, and antithetic sampling—VRPO achieves a significant stability improvement in aligning the diffusion model. [24] applys VRPO to LLaDA and fine-tune it with a preference model (reward model) on reasoning, math, code, and some other general alignment tasks.
IV.B.2 DIFFUSION-LLMs
@@ IV.B.3 DiffuGPT & DiffuLLaMA
IV.B.4 DREAM Series
DREAM 7B [7] is one of the most powerful open-source dLLMs to date. DREAM 7B achieves performance on par with, or exceeding, autoregressive models of similar size (e.g. it matches LLaMA3 8B [132] and Qwen 2.5-7B [133] on many benchmarks). A key to DREAM’s success is an optimized training recipe distilled from extensive experiments at smaller scales. [7] carefully explores the design choices on a 1B model and identify two especially valuable components: (1) AR weight initialization and (2) context adaptive noise scheduling.
@@ IV.B.4 DREAM Series
The Dream series has several follow-up works. Dream-Coder 7B [52] introduces the first fully open-source diffusion-based code model, adapted from Qwen2.5-Coder and trained on 322B tokens. It supports multiple decoding styles, including sketch-first generation, left-to-right decoding, and interleaved reasoning. DreamOn [51] resolves the fixed-canvas limitation of diffusion decoding by introducing special tokens ($|\text{expand}|>$, $|\text{delete}|>$) for variable-length generation. This innovation enables true flexible infilling and significantly improves performance on benchmarks such as HumanEval-Infilling.
IV.B.5 TESS 2
@@
IV.B.6 DiffuCoder
DiffuCoder [53] investigates the decoding behaviors of dLLMs through the lens of code generation. By training a 7B-parameter dLLM on a large code corpus, the authors reveal how dLLMs dynamically shift between autoregressive-like and parallel generation patterns. To stabilize preference optimization under diffusion training, they propose coupled-GRPO, a reinforcement learning algorithm that reduces variance via complementary mask sampling, yielding measurable improvements in code generation benchmarks.
IV.B.7 Seed Diffusion
Seed Diffusion [54] addresses the inductive biases and inefficiencies of discrete diffusion decoding. The authors propose a two-stage curriculum training strategy—beginning with mask-based corruption and transitioning to edit-based corruption—which strengthens the model’s self-correction ability. With further innovations in dynamic sequence expansion and contraction, Seed Diffusion establishes a new speed–quality Pareto frontier, reaching 2146 tokens/s while retaining strong generation accuracy.
IV.C Large Diffusion Multimodal Models
IV.C.1 Dimple
@@
[…]
IV.C.2 LaViDa
@@ IV.C.3 LLaDA-V
IV.D Large Unified Model
IV.D.1 MMaDA
@@ V.A.1 Low Corpus Utilization
V.A.2 Random Sampling of Time Index
@@ V.B.1 BERT Initialization
V.B.2 Autoregressive Model Initialization
@@
V.B.3 Autoregressive-then-Diffusion Training
@@
V.C Complementary Masking Technique
In [10], to ensure that all tokens participate in training, complementary masking constructs two complementary masked versions of each input sequence: $X_{t}$ and $X_{t}^{C}$. $X_{t}$ and $X_{t}^{C}$ have non-overlapping masked spans. For example, consider the sentence:
@@ V.C Complementary Masking Technique
V.D Masking Scheduling Technique
@@
V.D.1 Uniform Masking Scheduling
@@
The theoretical analyses of the optimal masking scheduling function remain scarce. [147] theoretically proves the optimality of cosine scheduling under the Fisher–Rao geometry.
V.D.2 Token-wise Masking Scheduling
Uniform masking scheduling uses the same scheduling for all tokens, which ignores the inherent variation in informativeness among different tokens. [42] introduces a token-wise masking schedule. The schedule defines a custom corruption probability $\alpha_{t}^{i}$ for each token position $i$ at timestep $t$, determined by:
@@ V.D.2 Token-wise Masking Scheduling
where $H(x_{0}^{i})$ denotes the entropy of the $i$-th token, measuring its information content, $S(t)$ is a sinusoidal scaling function that ensures zero informativeness contribution at $t=0$ and $t=T$, $\lambda$ is a hyperparameter controlling the strength of entropy influence.
@@
V.D.3 Blockwise SFT
For reasoning tasks, the prevailing approach adopts a blockwise semi-autoregressive decoding strategy. However, conventional supervised fine-tuning (SFT), applying random masking across the full response, does not reflect this procedure.
Let the response be partitioned into $M$ contiguous blocks $b^{(1)},\ldots,b^{(M)}$ of size $B$. For a given active block $a$, Blockwise SFT [58] defines:
  • Prefix $I^{(a)}_{\text{prefix}}$: tokens before block $a$, kept clean and fixed.
  • Active block $I^{(a)}$: tokens of block $a$, subject to stochastic masking.
  • Suffix $I^{(a)}_{\text{suffix}}$: tokens after block $a$, fully hidden.
In other words, the masking rule is:
$m_{i}=\begin{cases}0,&i\in I^{(a)}_{\text{prefix}}\quad\text{(clean prefix)},\\ \text{Bernoulli}(\pi),&i\in I^{(a)}\quad\text{(masked active block)},\\ 1,&i\in I^{(a)}_{\text{suffix}}\quad\text{(fully hidden suffix)},\end{cases}$ (78)
with $\pi\sim\text{Uniform}(10^{-3},1)$. Loss and gradient updates are computed only for the active block.
V.E Addressing Training-Testing Discrepancy